OpenStudy (anonymous):

Does the following make sense? f(x)=xsqrt(8-2x^2) , 0=

OpenStudy (anonymous):

Given \[f(x)=x \sqrt{8-2x^2} , 0\le x \le2\] for a graph  find surface area? i did it like that: \[f(x)=x \sqrt{8-2x^2} = \frac{ 1 }{ 2 } \int\limits_{0}^{2}2x \sqrt{8-2x^2}\] \[u=8-2x^2\] du=2xdx x=0 ---> u=8-2*0^2=8 x=2 ---> u=8-2*2^2=0 follows: \[\frac{ 1 }{ 2 }\int\limits_{8}^{0}\sqrt{u}du=\frac{ 1 }{ 2 }\int\limits_{8}^{0}u ^{\frac{ 1 }{ 2 }}du=\left[ \frac{ 1 }{ 3 } u ^{\frac{ 3 }{ 2 }}\right]_{8}^{0}=\frac{ 1 }{ 3 }*0-\frac{ 1 }{ 3 }*8^{\frac{ 3 }{ 2 }}=\frac{ 16\sqrt{2} }{ 3 }\] so answer is (16*2^1/2)/3?